why is it when a odd number is added to an odd number its always even, and when a odd number is multilied by a odd number its always odd.
Answer by RAY100(1637)

Lets just look at the visual representations
.
Addition,,,,0+0=00,,000+000=000000,,00000+00000=0000000000,, in short when we increment an odd number on top of an odd number ,,,it will be even.
.
If this was multiplication this would be an odd number taken an even number of times,,,(2)
.
.
.
Multiplication,,,,is repeated addition,,,but
.
1*1 = 0,,,,,
.
.
3*3 = 000
,,,,,,000
,,,,,,000,,,,,,,,,or the summation of an odd number by and odd number of sets.
.
.
5*5,,,,=,,,,00000
,,,,,,,,,,,,00000
,,,,,,,,,,,,00000
,,,,,,,,,,,,00000
,,,,,,,,,,,,00000,,,,,again an odd number taken an odd number of times
.
.
Hopefully this "adds" some insight... Please review several times.